OpenStudy (anonymous):

What is the remainder when 3^12 + 5^12 is divided by 13?

OpenStudy (shubhamsrg):

3^12 + 5^12 = 27^4 + 25^6 now 27 = 1 mod 13 =>27^4 = 1 mod 13 and 25 = -1 mod13 =>25^6 = 1 mod 13 bod leave remainder 1 each,,hence net remainder will be 1+1 or 2..

OpenStudy (anonymous):

or one can use fermat little theorem :)

OpenStudy (anonymous):

Fermat's little theorem states that if \(p\) is a prime number, then for any integer \(a\)\[a^{p-1} \equiv1 \ \ \ \text{mod} \ p\]

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Fermat's_little_theorem

OpenStudy (anonymous):

would you just show me that in this question?

OpenStudy (anonymous):

sure p=13 is a prime number so \[3^{13-1}=3^{12} \equiv1 \ \ \ \text{mod} \ 13\]\[5^{13-1}=5^{12} \equiv1 \ \ \ \text{mod} \ 13\]so\[3^{12}+5^{12} \equiv2 \ \ \ \text{mod} \ 13\]

OpenStudy (shubhamsrg):

learnt something new @mukushla ..thanks..

OpenStudy (anonymous):

np man :)